given f(x)=2(x)^2+8 and g(x)=5x-6 How can I find the function ( f+g)(x)?
Answer by jim_thompson5910(35256)

(f+g)(x) = f(x) + g(x)

(f+g)(x) = 2x^2+8 + 5x-6

(f+g)(x) = 2x^2 + 5x + 2